# Properties of observable mixed inertial/gravito-inertial modes in $\gamma$ Doradus stars

### v1

@@ D.1 Analytical solutions in the convective core
In this subsection, we describe small amplitude inertial oscillations in the convective core with no specified boundary conditions at the interface with the radiative zone. In an axially symmetric star, we search for solutions proportional to $e^{i(m\phi+\omega t)}$ where $m$ is the azimuthal number. The radial displacements and the Eulerian pressure perturbations are thus expressed as $\xi^{\rm conv}_{r}(r,\theta)e^{i(m\phi+\omega t)}$ and ${p^{\prime}}^{\rm conv}(r,\theta)e^{i(m\phi+\omega t)}$ respectively. In an isentropic convective core of uniform density, perturbations of low frequency are governed by the Poincaré equation $\Delta\Psi=s^{2}\frac{\partial\Psi}{\partial z}$ where $\Psi$ is proportional to ${p^{\prime}}^{\rm conv}$ (Wu 2005). In a full sphere, separable solutions are obtained using the ellipsoidal coordinates ($x_{1},x_{2},\phi$) defined in terms of the Cartesian coordinates as:
@@
where $\mu_{s}=1/s$, $x_{1}\in[\mu_{s},1]$ and $x_{2}\in[-\mu_{s},\mu_{s}]$. With these coordinates, the surface of the sphere is given by $\{x_{1}\in[\mu_{s},1],x_{2}=\mu_{s}\}\cup\{x_{1}=\mu_{s},x_{2}\in[-\mu_{s},\mu_{s}]\}$.
@@
These solutions oscillate freely at the convective/radiative interface because no boundary condition is imposed there. If the $\xi^{\rm conv}_{r}(r=r_{c})=0$ boundary condition is enforced, one recovers the eigenfrequencies of pure inertial modes presented in Appendix A which satisfy Eq. 22 or equivalently $\gamma_{\ell_{i}}^{m}(s)=0$.
@@ D.2 Analytical solutions in the radiative zone
With the Cowling approximation, the equations governing the radial part of the oscillation, $\zeta$ and $p^{\prime}$, are:
@@
D.3 Mixed inertial/gravito-inertial modes
To produce mixed inertial/gravito-inertial modes, the inertial and gravito-inertial oscillations have to match at the convective/radiative interface in the sense that $\xi_{r}$ and $\delta p$ have to be continuous there. However, as illustrated in Fig. 17 for the $(\ell=1,m=-1)$ vs $(\ell_{i}=3,m=-1)$ resonance, these matching conditions can not be fulfilled with only one separable solutions in each domain. We thus need more general solutions on both sides of the interface. This can be done by considering linear combinations of the separable solutions derived in the two previous sections. Namely, in the radiative zone :
@@ D.3 Mixed inertial/gravito-inertial modes
At the convective core/radiative zone interface, the density is continuous while its radial derivative is discontinuous. In such a case, the continuity condition on $\xi_{r}$ and $\delta p$ is equivalent to the continuity of $\xi_{r}$ and $p^{\prime}$, thus to ${\xi}^{\rm conv}_{r}(r=r_{c}^{-})={\xi}^{\rm rad}_{r}(r=r_{c}^{+})$ and $p^{\prime\rm conv}(r=r_{c}^{-})=p^{\prime\rm rad}(r=r_{c}^{+})$.
@@
To get approximate analytical solutions, we may try to truncate this infinite system of equations to a finite system. We first consider the case of the resonance between the Kelvin modes $(\ell=1,m=-1)$ and the $(\ell_{i}=3,m=-1)$ inertial mode near $s_{*}=11.32$. To analyze the geometrical matching between these modes at $r=r_{c}$, the upper panel of Fig. 39 shows how $H_{0}^{-1}(\theta,s_{*})$, projects onto Legendre polynomials with same $m$ and equatorial parity. We observe that the two main contributions come from $P_{1}^{-1}$ and $P_{3}^{-1}$. Conversely, the lower panel of Fig. 39 shows that the inertial mode $P_{3}^{-1}$ mostly projects onto $H_{0}^{-1}(s_{*})$ and $H_{-2}^{-1}(s_{*})$. This suggests to truncate the infinite system to linear combinations of $P_{3}^{-1}$ and $P_{1}^{-1}$ on the convective side and of $H_{0}^{-1}(s)$ and $H_{-2}^{-1}(s)$ on the radiative side. For consistency one should also consider the modes excited by $H_{-2}^{-1}(s_{*})$ and $P_{1}^{-1}$ at the interface. The projections of $H_{-2}^{-1}(s)$ and $P_{1}^{-1}$ onto Legendre polynomials and Hough functions respectively are displayed in Fig. 40. It shows that, except for the contribution of $P_{5}^{-1}$ to $H_{-2}^{-1}(s)$, the two functions $H_{-2}^{-1}(s)$ and $P_{1}^{-1}$ are dominantly projected onto respectively $P_{3}^{-1}$ and $P_{1}^{-1}$, and $H_{0}^{-1}(s_{*})$ and $H_{-2}^{-1}(s)$. We thus expect that the truncation of the infinite system to $P_{1}^{-1}$, $P_{3}^{-1}$, $H_{0}^{-1}(s)$ and $H_{-2}^{-1}(s)$ provides an approximate representation of the mixed inertial/gravito-inertial modes near $s_{*}=11.32$.
Figure 39: (Upper panel) Projection of the Hough function $H_{k=0}^{m=-1}(s_{*}=11.32)$ on the associated Legendre polynomial of the same $m$ and parity. (Lower panel) Projection of the associated Legendre polynomial $P_{\ell_{i}=3}^{m=-1}$ on the Hough functions of the same $m$ and parity determined at $s_{*}=11.32$.
Figure 40: (Upper panel) Projection of the Hough function $H_{k=-2}^{m=-1}(s_{*}=11.32)$ on the associated Legendre polynomial of the same $m$ and parity. (Lower panel) Projection of the associated Legendre polynomial $P_{\ell_{i}=1}^{m=-1}$ on the Hough functions of the same $m$ and parity determined at $s_{*}=11.32$.
@@
Having a non trivial solution for this system requires that its determinant vanishes. This can be written as:
@@
We can simplify this equation using the fact that $\Omega/N_{0}\ll 1$ in $\gamma$ Doradus stars. However, one should pay attention that $\gamma_{3}$ is also small as we consider spin parameters close to $s_{*}$ and $\gamma_{3}(s_{*})=0$. We obtain:
@@
Close to $s_{*}$ we can write $\gamma_{3}\approx\gamma^{\prime}_{3}(s-s_{*})$, while the other terms in ${\cal A}$ et ${\cal B}$ remain approximately equal to their value at $s_{*}$. Keeping the dominant terms, we get:
@@
To conclude, we note that up to the (infinite) system of equations Eq. 55 describing the continuity conditions, the present model is identical to that of Tokuno & Takata (2022). From there, we improved the continuity conditions by adding the contribution of the Hough function $H_{k-2}^{m}$ on the radiative side and of the Legendre polynomial $P_{\ell_{i}-2}^{m}$ mode on the convective side. While we argue this is enough to approximate the resonances considered in this paper, we stress that this model cannot be generalised to all dips. Indeed, for other resonances (e.g. the $(\ell=2,m=0)$ vs $(\ell_{i}=4,m=0)$ resonance), the projections of the $H_{k}^{m}$ Hough function on the Legendre polynomials and of the $P_{\ell_{i}}^{m}$ Legendre polynomial on the Hough functions strongly suggest that many more free oscillation modes will be involved leading to mixed modes much more complex than the ones considered here.

### v2

@@ D.1 Analytical solutions in the convective core
In this subsection, we describe small amplitude inertial oscillations in the convective core with no specified boundary conditions at the interface with the radiative zone. In an axially symmetric star, we search for solutions proportional to $e^{i(m\phi+\omega t)}$ where $m$ is the azimuthal number. The radial displacement and the Eulerian pressure perturbations are thus expressed as $\xi^{\rm conv}_{r}(r,\theta)e^{i(m\phi+\omega t)}$ and ${p^{\prime}}^{\rm conv}(r,\theta)e^{i(m\phi+\omega t)}$ respectively. In an isentropic convective core of uniform density, perturbations of low frequency are governed by the Poincaré equation $\Delta\Psi=s^{2}\frac{\partial\Psi}{\partial z}$ where $\Psi$ is proportional to ${p^{\prime}}^{\rm conv}$ (Wu 2005). In a full sphere, separable solutions are obtained using the ellipsoidal coordinates ($x_{1},x_{2},\phi$) defined in terms of the Cartesian coordinates as
@@
where $\mu_{s}=1/s$, $x_{1}\in[\mu_{s},1]$, and $x_{2}\in[-\mu_{s},\mu_{s}]$. With these coordinates, the surface of the sphere is given by $\{x_{1}\in[\mu_{s},1],x_{2}=\mu_{s}\}\cup\{x_{1}=\mu_{s},x_{2}\in[-\mu_{s},\mu_{s}]\}$.
@@
These solutions oscillate freely at the convective-radiative interface because no boundary condition is imposed there. If the $\xi^{\rm conv}_{r}(r=r_{c})=0$ boundary condition is enforced, one recovers the eigenfrequencies of pure inertial modes presented in Appendix A which satisfy Eq. 22 or equivalently $\gamma_{\ell_{i}}^{m}(s)=0$.
@@ D.2 Analytical solutions in the radiative zone
Within the Cowling approximation, the equations governing the radial part of the oscillation, $\zeta$ and $p^{\prime}$, are:
@@
D.3 Modes of mixed inertial and gravito-inertial nature
To produce mixed modes, the inertial and gravito-inertial oscillations have to match at the convective-radiative interface in the sense that $\xi_{r}$ and $\delta p$ have to be continuous there. However, as illustrated in Fig. 17 for the $(\ell=1,m=-1)$ vs $(\ell_{i}=3,m=-1)$ resonance, these matching conditions can not be fulfilled with only one separable solutions in each domain. We thus need more general solutions on both sides of the interface. This can be done by considering linear combinations of the separable solutions derived in the two previous sections. Namely, in the radiative zone :
@@ D.3 Modes of mixed inertial and gravito-inertial nature
At the interface between the convective core and the radiative zone, the density is continuous while its radial derivative is discontinuous. In such a case, the continuity condition on $\xi_{r}$ and $\delta p$ is equivalent to the continuity of $\xi_{r}$ and $p^{\prime}$, thus to ${\xi}^{\rm conv}_{r}(r=r_{c}^{-})={\xi}^{\rm rad}_{r}(r=r_{c}^{+})$ and $p^{\prime\rm conv}(r=r_{c}^{-})=p^{\prime\rm rad}(r=r_{c}^{+})$.
@@
To get approximate analytical solutions, we may try to truncate this infinite system of equations to a finite system. We first consider the case of the resonance between the Kelvin modes $(\ell=1,m=-1)$ and the $(\ell_{i}=3,m=-1)$ inertial mode near $s_{*}=11.32$. To analyse the geometrical matching between these modes at $r=r_{c}$, the upper panel of Fig. 39 shows how $H_{0}^{-1}(\theta,s_{*})$, projects onto Legendre polynomials of same $m$ and equatorial parity. We observe that the two main contributions come from $P_{1}^{-1}$ and $P_{3}^{-1}$. Conversely, the lower panel of Fig. 39 shows that the inertial mode $P_{3}^{-1}$ mostly projects onto $H_{0}^{-1}(s_{*})$ and $H_{-2}^{-1}(s_{*})$. This suggests to truncate the infinite system to linear combinations of $P_{3}^{-1}$ and $P_{1}^{-1}$ on the convective side and of $H_{0}^{-1}(s)$ and $H_{-2}^{-1}(s)$ on the radiative side. For consistency one should also consider the modes that are excited by $H_{-2}^{-1}(s_{*})$ and $P_{1}^{-1}$ at the interface. The projections of $H_{-2}^{-1}(s)$ and $P_{1}^{-1}$ onto Legendre polynomials and Hough functions respectively are displayed in Fig. 40. It shows that, except for the contribution of $P_{5}^{-1}$ to $H_{-2}^{-1}(s)$, the two functions $H_{-2}^{-1}(s)$ and $P_{1}^{-1}$ are dominantly projected onto respectively $P_{3}^{-1}$ and $P_{1}^{-1}$, and $H_{0}^{-1}(s_{*})$ and $H_{-2}^{-1}(s)$. We thus expect that the truncation of the infinite system to $P_{1}^{-1}$, $P_{3}^{-1}$, $H_{0}^{-1}(s)$, and $H_{-2}^{-1}(s)$ provides an approximate representation of the mixed modes near $s_{*}=11.32$.
Figure 39: Upper panel: Projection of the Hough function $H_{k=0}^{m=-1}(s_{*}=11.32)$ on the associated Legendre polynomial of the same $m$ and parity. Lower panel: Projection of the associated Legendre polynomial $P_{\ell_{i}=3}^{m=-1}$ on the Hough functions of the same $m$ and parity determined at $s_{*}=11.32$.
Figure 40: Upper panel: Projection of the Hough function $H_{k=-2}^{m=-1}(s_{*}=11.32)$ on the associated Legendre polynomial of the same $m$ and parity. Lower panel: Projection of the associated Legendre polynomial $P_{\ell_{i}=1}^{m=-1}$ on the Hough functions of the same $m$ and parity determined at $s_{*}=11.32$.
@@
Having a non-trivial solution for this system requires that its determinant vanishes. This can be written as:
@@
We can simplify this equation using the fact that $\Omega/N_{0}\ll 1$ in $\gamma$ Dor stars. However, one should pay attention that $\gamma_{3}$ is also small as we consider spin parameters close to $s_{*}$ and $\gamma_{3}(s_{*})=0$. We obtain:
@@
Close to $s_{*}$ we can write $\gamma_{3}\approx\gamma^{\prime}_{3}(s-s_{*})$, while the other terms in ${\cal A}$ and ${\cal B}$ remain approximately equal to their value at $s_{*}$. Keeping the dominant terms, we get:
@@
To conclude, we note that up to the (infinite) system of equations Eq. 55 describing the continuity conditions, the present model is identical to that of Tokuno & Takata (2022). From there, we improved the continuity conditions by adding the contribution of the Hough function $H_{k-2}^{m}$ on the radiative side and of the Legendre polynomial $P_{\ell_{i}-2}^{m}$ mode on the convective side. While we argue this is enough to approximate the resonances considered in this paper, we stress that this model cannot be generalised to all dips. Indeed, for other resonances (e.g. the $(\ell=2,m=0)$ vs $(\ell_{i}=4,m=0)$ resonance), the projections of the $H_{k}^{m}$ Hough function on the Legendre polynomials and of the $P_{\ell_{i}}^{m}$ Legendre polynomial on the Hough functions strongly suggest that many more free oscillation modes will be involved leading to much more complex mixed modes than the ones considered here.